OpenStudy (sosa4954):

2/3x + 90 = 80 would x=-10

OpenStudy (anonymous):

2/3 x +90=80 2/3 x = -10 x= -10(3/2) x=-15

OpenStudy (leader):

\[\frac{ 2 }{ 3 }x+90=80~would~x=10\]

OpenStudy (leader):

Is that the question.

OpenStudy (leader):

@Sosa4954

OpenStudy (sosa4954):

i just got a different answer i got x=-30/2

OpenStudy (leader):

x=-15

OpenStudy (sosa4954):

the question is \[\frac{ 2 }{ 3 }x + 90=80\]

OpenStudy (leader):

2/3 x+90=80 Subtract 90. 2/3x=-10 Multiply by 3. 2x=-30 Divide by 2. x=-15

OpenStudy (sosa4954):

u guys did it so differently i did -90 on both sides and that left me with 2/3x=-10 than i had to cross multiply so i had to multiply 3/2 to get x alone and i did it on both sides so it left me with x = -30/2

OpenStudy (anonymous):

-30/2 simplifies to -15, so you're right too.

OpenStudy (sosa4954):

oh ok thanks lol